Two long parallel wires carry currents of 3.59 A and 4.41 A. The magnitude of the force per unit lengh acting on each wire is 6.39
Answer #1

F=IBL

B=U0*I/(2*pi*r)

6.39*10^-5 =(3.59*4.41*2*10^-7 )/r

r=20m
Answer #2

dF/dl=munot)I1I2/2pi)d

dF/dl=2*10^-7*3.59*4.41/d=6.39*10^-5

d=0.049 m=4.9cm (ans)